# FireFly-S: Exploiting Dual-Side Sparsity for Spiking Neural Networks Acceleration with Reconfigurable Spatial Architecture

### v2

@@ V-A Dataflow Orchestrator
As discussed in Section IV, spatial architectures require a consistent dataflow pattern. However, achieving this consistency presents a challenge due to the conflicting dataflow requirements between the output of one layer and the input of the next. Specifically, the membrane potential storage-free calculation in FireFly-S necessitates placing the output channel loop outside the temporal loop. This arrangement contrasts with the input dataflow of the subsequent layer, which requires the channel loop to be inside the temporal loop for efficient processing. Furthermore, convolution operations introduce an additional complexity: the need for the im2col operation to transform the input feature map into a matrix format suitable for convolution.
[…]
@@
[…]
@@ V-B Dual-side Sparsity Detector
The next stage utilizes the one-hot vector $y$ to generate a new vector with bits set from the least significant bit up to the index where ’1’ is set in $y$. For example, if $y$ is 4’b0100, which indicates the third position, the generated vector would be 4’b0111. This vector is then subjected to a bitwise AND operation with the weight mask to pinpoint active indices that align with the mask. Continuing with the example, if the weight mask is 4’b1101, the result of the AND operation between 4’b0111 and 4’b1101 would be 4’b0101. This result, obtained at CLK3, allows the detector at the following cycle (CLK4) to determine the precise storage offset — 2 in this example — of the weight data corresponding to the ’1’ in the original one-hot vector 4’b0100 using an adder tree. At CLK5, the calculated offset is accumulated into an offset register. This accumulated offset is then added to the base address stored in a base address register to generate the final address for accessing the weight data, thus preparing the read command for the specific weight data. Finally, at CLK6, the weight data is fetched from the RAM and used to peform IF/LIF neuron model computation. It is important to note that the membrane potential is stored in a register, and due to the dedicated dataflow, the voltage for each time step is temporally stored in the register and reused for the next time step.
For clarity, Fig. 7 does not explicitly illustrate the interactions with bias and threshold RAMs as depicted in Fig. 5. However, these interactions are straightforward: each output channel’s neuron dynamics require the addition of bias, and the membrane potential must be compared with the threshold to decide if a spike should be generated. Since both bias and threshold are assigned per channel (reflecting per-channel quantization), and because the bias is always added and the threshold is always checked against the membrane potential to determine spike generation, the reading addresses for both are represented by an incrementing counter.
Moreover, to improve efficiency during the dual-side sparse acceleration process involving weight and spike vectors, a specific optimization is employed. Typically, if the bitwise AND operation between the spike vectors and the weight vector bitmap results in all zeros, it would lead to a wasted cycle due to the absence of valid nonzero pairs. To counter this inefficiency, we incorporate the bias, treated as a constant input of ’1’ with a corresponding weight equal to the bias value, which is processed in each time step. By designing logic to absorb the cycle that would have been wasted on detection into the cycle used for adding the bias, we ensure that the bias effectively replaces the ineffective cycle. This eliminates unnecessary idle clock cycles and enhances the efficiency of the Dual-side Sparsity Detector.
@@ VI-B Parallelism Setup
Model LUTs FFs CARRY8s BRAMs
@@
Module LUTs FFs CARRY8s BRAMs
@@
Workload per $\text{Workload per Unit}=\text{MAC}\times\frac{s_{1}\times s_{2}}{P_{C_{o}}}.$ (10)
@@ VI-C Resource and Performance Evaluation
    SyncNN calculates power usage by measuring the active power consumption of the FPGA board (24.5 W) and subtracting the static power measured when the board is idle (24.1 W). Conversely, it is believed that most existing research acquires power metrics directly from the Vivado report.
@@
Table IV details the resource utilization of FireFly-S across various model deployments on the Xilinx Ultrascale KV260, equipped with an xczu5ev chip. As model sizes increase, we observe a corresponding rise in the utilization of the resources, this trend is attributable to an increase in the number of layers and the characteristic of spatial architectures. Notably, the efficient 4-bit quantization and the intrinsic spike-based operation of the SNNs allow inference processes to operate without any DSP Slices, relying solely on LUTs. Additionally, the application of model size compression through 4-bit quantization and high sparsity levels significantly reduces BRAM usage. For example, espite the 8-bit quantized weight parameters of SCNN9 still requiring 9.5Mb—surpassing the chip’s BRAM capacity of 5.1Mb and 3.5Mb of Distributed RAM—the reduction techniques lower the BRAM demand to just 53%. This enhanced efficiency supports the feasible deployment of larger models on the chip.
@@
    Values are estimates based on the Fig. 9 of SyncNN[8], as no exact values were provided in the original text.
@@
    The symbol ”$\text{X}_{\text{@Y}}$” indicates that accuracy = Xand sparsity = Y%.
Finally, it is important to highlight that although our chosen device, the xczu5ev, is an edge device, FireFly-S still demonstrates significant improvements across all benchmarks. By utilizing a larger device such as the xczu7ev in the ZCU104, the inference performance can be further enhanced due to the substantially greater number of LUT resources available, which alleviates the bottlenecks experienced with the xczu5ev and supports higher levels of parallelism. Additionally, it is noteworthy that our model is fully mapped on-chip, which means that each layer utilizes fewer resources compared to an overlay architecture. Despite this, our design achieves remarkable performance without any off-chip memory access, thereby eliminating external power consumption typically associated with overlay architectures.
@@ VII Conclusions
In this work, we introduce a high-energy-efficient and reconfigurable hardware accelerator for SNNs, leveraging a software-hardware co-design approach to optimize energy efficiency during SNN inference. On the software side, our approach includes a quantization-aware and pruning-aware training method, which enables the training of SNN models at 4-bit quantization and 85-95% sparsity with competitive accuracy loss. This training strategy not only allows the deployment of large SNN models on-chip but also enhances calculation logic utilization by introducing weight sparsity through pruning. On the hardware side, we capitalize on the natural sparsity of spikes inherent to SNNs and the induced sparsity of pruned weights to accelerate processing. This is effectively exploited by our Dual-side Sparsity Detector, which accelerates processing by leveraging both types of sparsity and skipping the unnecessary calculations. Additionally, our architecture is designed to be fully pipelined, enhancing throughput. The Dataflow Orchestrator plays a critical role in maintaining consistent dataflow between layers, minimizing stalls, and simplifying logic to reduce resource consumption. We validate the reconfigurability of our FireFly-S accelerator by evaluating various deep SNN models across multiple datasets. Utilizing commercially available FPGA edge devices offers a more practical and feasible solution compared to other existing hardware platforms. Future work will focus on enhancing the SNN hardware infrastructure further, aiming for ultra-high compression rates and addressing efficiency declines at extreme sparsity levels in Bitmap representation.

### v3

@@ V-A Dataflow Orchestrator
As discussed in Section IV, spatial architectures require a consistent dataflow pattern. However, achieving this consistency presents a challenge due to the conflicting dataflow requirements between the output of one layer and the input of the next layer. Specifically, the membrane potential storage-free calculation in FireFly-S necessitates placing the output channel loop outside the temporal loop. This arrangement contrasts with the input dataflow of the subsequent layer, which requires the channel loop to be inside the temporal loop for efficient processing. Furthermore, convolution operations introduce an additional complexity: the need for the im2col operation to transform the input feature map into a matrix format suitable for convolution.
Figure 6: Dataflow Orchestrator in FireFly-S, illustrated with an example where $(T,C_{o},C_{i},F_{w_{i}},F_{h_{i}},K_{w_{i}},K_{h_{i}})=(2,\ 2,\ 3,\ 4,\ 4,\ 3,\ 3)$ and $(P_{C_{o}},P_{C_{i}})=(1,\ 1)$. The ”Priority” column in the table indicates the order in which stride is taken, with higher numbers representing higher priority.
@@
Figure 7: Schematic of the Dual-side Sparsity Detector in FireFly-S, illustrated with configuration parameters $(C_{o},C_{i},K_{w_{i}},K_{h_{i}})=(4,\ 8,\ 3,\ 3)$ and parallelism parameters $(P_{C_{o}},P_{C_{i}})=(2,\ 4)$.
@@ V-B Dual-side Sparsity Detector
The next stage utilizes the one-hot vector $y$ to generate a new vector with bits set from the least significant bit up to the index where ’1’ is set in $y$. For example, if $y$ is 4’b0100, which indicates the third position, the generated vector would be 4’b0111. This vector is then subjected to a bitwise AND operation with the weight mask to pinpoint active indices that align with the mask. Continuing with the example, if the weight mask is 4’b1101, the result of the AND operation between 4’b0111 and 4’b1101 would be 4’b0101. This result, obtained at CLK3, allows the detector at the following cycle (CLK4) to determine the precise storage offset — 2 in this example — of the weight data corresponding to the ’1’ in the original one-hot vector 4’b0100 using an adder tree. At CLK5, the calculated offset is accumulated into an offset register. This accumulated offset is then added to the base address stored in a base address register to generate the final address for accessing the weight data, thus preparing the read command for the specific weight data. Finally, at CLK6, the weight data is fetched from the RAM and used to perform IF/LIF neuron model computation. It is important to note that the membrane potential is stored in a register, and due to the dedicated dataflow, the voltage for each time step is temporally stored in the register and reused for the next time step.
For clarity, Fig. 7 does not explicitly illustrate the interactions with bias and threshold RAMs as depicted in Fig. 5. However, these interactions are straightforward: each output channel’s neuron dynamics require the addition of bias, and the membrane potential must be compared with the threshold to determine whether a spike should be generated. Since both bias and threshold are assigned per channel (reflecting per-channel quantization), and because the bias is always added and the threshold is always checked against the membrane potential to determine spike generation, the reading addresses for both are represented by an incrementing counter.
Moreover, to improve efficiency during the dual-side sparse acceleration process involving weight and spike vectors, a specific optimization is employed. Typically, if the bitwise AND operation between the spike vectors and the weight vector bitmap results in all zeros, it would lead to a wasted cycle due to the absence of valid non-zero pairs. To counter this inefficiency, we incorporate the bias, treated as a constant input of ’1’ with a corresponding weight equal to the bias value, which is processed in each time step. By designing logic to absorb the cycle that would have been wasted on detection into the cycle used for adding the bias, we ensure that the bias effectively replaces the ineffective cycle. This eliminates unnecessary idle clock cycles and enhances the efficiency of the Dual-side Sparsity Detector.
@@ VI-B Parallelism Setup
Model LUTs $F$s CARRY8s BRAMs
@@
Module LUTs $F$s CARRY8s BRAMs
@@
$\text{Workload per Unit}=\text{MAC}\times\frac{s_{1}\times s_{2}}{P_{C_{o}}}.$ (10)
@@ VI-C Resource and Performance Evaluation
    SyncNN calculates power usage by measuring the active power consumption of the FPGA board (24.5 W) and subtracting the static power measured when the board is idle (24.1 W). By contrast, it is believed that most existing research acquires power metrics directly from the Vivado report.
@@
Table IV details the resource utilization of FireFly-S across various model deployments on the Xilinx Ultrascale KV260, equipped with an xczu5ev chip. As model sizes increase, we observe a corresponding rise in the utilization of the resources. This trend is attributable to an increase in the number of layers and the characteristics of spatial architectures. Notably, the efficient 4-bit quantization and the intrinsic spike-based operation of the SNNs allow inference processes to operate without any DSP Slices, relying solely on LUTs. Additionally, the application of model size compression through 4-bit quantization and high sparsity levels significantly reduces BRAM usage. For example, although the 8-bit quantized weight parameters of SCNN9 still require 9.5Mb—surpassing the chip’s BRAM capacity of 5.1Mb and 3.5Mb of Distributed RAM—the reduction techniques lower the BRAM demand to just 53%. This enhanced efficiency supports the feasible deployment of larger models on the chip.
@@
    Values are estimates based on Fig. 9 of SyncNN[8], as no exact values were provided in the original text.
@@
    The symbol $\text{X}_{\text{@Y}}$” indicates that accuracy = X% and sparsity = Y%.
Notably, despite the resource constraints of the edge-tier xczu5ev device, FireFly-S still demonstrates significant improvements across all benchmarks. By utilizing a larger device such as the xczu7ev in the ZCU104, the inference performance can be further enhanced due to the substantially greater number of LUT resources available, which alleviates the bottlenecks experienced with the xczu5ev and supports higher levels of parallelism. Additionally, it is noteworthy that our model is fully mapped on-chip, which means that each layer utilizes fewer resources compared to an overlay architecture. Despite this, our design achieves remarkable performance without any off-chip memory access, thereby eliminating external power consumption typically associated with overlay architectures.
@@ VII Conclusions
In this work, we introduce a highly energy-efficient and reconfigurable hardware accelerator for SNNs, leveraging a software-hardware co-design approach to optimize energy efficiency during SNN inference. On the software side, our approach includes a quantization-aware and pruning-aware training method, which enables the training of SNN models at 4-bit quantization and 85–95% sparsity with competitive accuracy loss. This training strategy not only allows the deployment of large SNN models on-chip but also enhances calculation logic utilization by introducing weight sparsity through pruning. On the hardware side, we capitalize on the natural sparsity of spikes inherent to SNNs and the induced sparsity of pruned weights to accelerate processing. This is effectively exploited by our Dual-side Sparsity Detector, which accelerates processing by leveraging both types of sparsity and skipping the unnecessary calculations. Additionally, our architecture is designed to be fully pipelined, enhancing throughput. The Dataflow Orchestrator plays a critical role in maintaining consistent dataflow between layers, minimizing stalls, and simplifying logic to reduce resource consumption. We validate the reconfigurability of our FireFly-S accelerator by evaluating various deep SNN models across multiple datasets. Utilizing commercially available FPGA edge devices offers a more practical and feasible solution compared to other existing hardware platforms. Future work will focus on enhancing the SNN hardware infrastructure further, aiming for ultra-high compression rates and addressing efficiency declines at extreme sparsity levels in Bitmap representation.